# From Friendship Networks to Classroom Dynamics: Leveraging Neural Networks, Instrumental Variable and Genetic Algorithms for Optimal Educational Outcomes

### v4

@@ 5.2 Positive peer effect from ‘good students’
We estimate Eq(3) and confirm the presence of positive peer effect from ‘good students’ (students who did well in elementary school). In particular, if student $i$’s friends’ weighted average 6th-grade class quantile improves by 10%, his 8th-grade cognitive ability test score improves by 0.1082 on a scale of 5, or equivalently, 13% of standard deviation. To put it more concretely, the average cognitive test score is 0.3608 which lands the student at the 43rd quantile of the score distribution (the score distribution is left-skewed). An 0.1082 increase in cognitive test score improves the student’s ranking from 43rd quantile to 51st quantile. In Table 2, we show four specifications’ results.
  • Linear-in-means model without classroom random effect
  • Linear-in-means model with classroom random effect
  • Instrumental variable without classroom random effect
  • Instrumental variable with classroom random effect
@@
Linear-in-means Linear-in-means IV IV
Peer’s Rank 0.973∗∗∗ 0.980∗∗ 1.075∗∗∗ 1.082∗∗
@@
Age -0.012∗∗∗ -0.012∗∗∗ -0.012∗∗∗ -0.012∗∗∗
@@
Sex -0.010 -0.010 -0.011 -0.010
(0.017) (0.017) (0.017) (0.017)
Father’s education 0.020∗∗∗ 0.017∗∗∗ 0.020∗∗∗ 0.018∗∗∗
(0.006) (0.006) (0.006) (0.006)
Mother’s education 0.006 0.006 0.005 0.005
(0.006) (0.006) (0.006) (0.006)
Ethnic nationality 0.013 0.022 0.012 0.021
(0.040) (0.040) (0.040) (0.040)
Number of observations 5,860 5,860 5,860 5,860
Adjusted $R^{2}$ 0.382 0.382
- Log Likelihood 5,845.390 5,845.291
Results from first stage
Peer’s Rank 0.905∗∗∗ 0.905∗∗∗
(0.015) (0.015)
Number of observations 5,860 5,860
Adjusted $R^{2}$ 0.755 0.755
@@
We also implement more standard methods (Baltagi, 1981; Balestra and Varadharajan-Krishnakumar, 1987) for IV with random effect specification to estimate Eq(3). The estimated coefficients for peer effect are 0.975 (0.268) and 1.075 (0.270), respectively. The numbers in the parenthesis are standard errors. The regression results for all four specifications and all estimation methods are very close to each other, demonstrating robustness. See source code in supplemental material for details.
The regression results opens up the strategic scope of classroom assignments. Making good students the central nodes of the social network increases the entire classroom’s peer effect. On the other hand, if the estimated $\hat{\beta}$ were significantly negative, we would aim to make good students peripheral nodes. The sign of the estimated peer effect determines two opposing sets of policies.
@@ 5.3 Optimal classroom assignment policy design
Under the raw assignment, many students are negatively influenced by two disruptive peers as indicated by the two darker rows and columns of the Raw Q2 plot in Figure 4. GA separates many pairs of students who suffer from such large negative peer effects by ensuring that these two disruptive peers are popular among fewer students as evidenced by the reduced number of dark grids in GA Q1 and GA Q2 relative to Raw Q1 and Raw Q2. Moreover, GA Q2 shows that the entire classroom mostly consists of positively influencing friend pairs. Consequently, students’ average predicted peer effect, Eq(5), increases by 1.9% under GA assignment policy. Unfortunately, maximizing the average peer effect inadvertently leads to severe negative peer effects for some students. In GA Q2, many students face extremely negative peer effects.
@@
We deploy AFGA to address this inequity issue. AF-GA Q1 and AF-GA Q2 in Figure 4 show a much more equitable educational outcome. None of the students suffer from extremely negative peer effects as in the GA Q2. However, though it is not clear in Figure 4, the average peer effect for AFGA (compared to raw assignment, the improvement is 1.2%) is lower than GA (1.9%), which is expected since the objective function of AFGA Eq(6) does take into account equity. If the school principal believes this particular policy designed by AFGA overvalues equity and sacrifices efficiency too much, he can decrease $\phi$ and $\rho$ for a more efficient outcome, however, the outcome will be less equitable. Note that the school principal can predict peer effect, $\tilde{\Omega}$, for all students before implementing the policy, hence he can always choose $\phi$ and $\rho$ based on his preference over efficiency versus equity.
@@ 6 Conclusion
Broader impact: This three-step empirical framework is extendable to other group assignment challenges where peer effects play a pivotal role. Some of the applicable contexts, in which peer effects have been empirically proven crucial and group assignments are common practices, are workplace teams (Anderson et al., 2008; Bandiera et al., 2010; Mas and Moretti, 2009), healthcare interventions (Pianese and Belfiore, 2021; Pagkas-Bather et al., 2020) and community programs (Valente et al., 2003; Yu et al., 2024; Cero et al., 2024). Similar to classroom assignments, navigating the trade-off between efficiency and equity in these situations can present moral dilemmas.
Limitations and future work: The scope of this paper does not include the theoretical properties of the three-step empirical framework. Some of the future research directions are (1) characterizing the conditions under which the friendship formation model is identifiable, (2) proving consistency of the IV approach when $\Omega$ is not taken as the ground truth, and (3) showing the error bound of GA as the number of search goes to infinity.
Appendix A Further literature review
A.1 Comparison between our model and other causal inference methods under networked interference
Ma and Tresp (2021); Jiang and Sun (2022); Huang et al. (2023) all adapt GNN for causal inference. The core problem in treatment effect estimation is that the treated group and control group have imbalanced features, both their individual feature as well as their neighborhood information.
For example, students who enroll themselves in a tutoring program are inherently different (possibly more self-motivating) from those who don’t. At the same time, these tutoring program participants may have friends who tend to do well academically. Hence, if we observe tutoring program participants on average do better than other students, we cannot make the casual statement that attending the tutoring program improves students’ grades. Their better performance could be due to their self-motivation or more helpful friends.
To resolve these two imbalanced data problems, all three papers encode their raw data into latent feature space. They formalize mathematically the concept of balancing the distributions of latent features for treated and control groups as part of the loss function. Ma and Tresp (2021); Huang et al. (2023) both directly penalize some distance measures of the two distributions (latent feature distribution of the treated group and that of the control group). Jiang and Sun (2022), in contrast, trains a discriminator that learns about the probability of treatment based on those latent features and eventually outputs a set of latent features such that the discriminator can no longer distinguish the treatment probability between the treated group and control group. For an imbalanced network problem, all three papers also collapse all neighbors’ treatment information into a scalar, percentage of neighbors being treated. After the transformation of neighbor information, they address the network imbalance problems just like how they resolve the treatment imbalance.
Additionally, Huang et al. (2023) controls for unobserved confounding by adding the treatment prediction loss to its loss function. The weights for balancing the distributions and treatment prediction are tuning parameters in their case.
The challenge faced by those three papers is rooted in nonrandom treatment. In our case, the class assignment itself is random for 5860 students and 139 classes. Nevertheless, causal inference is not a straightforward linear regression because friendship formation is not random. We exploit the random assignment to generate instrumental variable for friends’ weighted average 6th-grade class rank. By standard two-stage least-square estimation, we can consistently estimate the parameter of interest $\beta$ in Equation (3). The instrumental variable identification strategy solves both the ‘treatment’ imbalance (we can perceive friends’ weighted average 6th-grade class quantile as continuous treatment) and unobserved confounding problems.
Next, we compare the policy optimization method of our paper and Ma and Tresp (2021). Ma and Tresp (2021) build a neural network to solve for optimal policy. This is doable in their case as the network is predetermined and treatment assignment does not change the network. In contrast, our class assignment changes the network structure. For example, if students $i$ and $j$ are not in the same class, we assume that there is zero probability of being friends, on the other hand, when they are in the same classroom, there is always a nonzero probability of them being friends due to the use of softmax function. Network structure is assignment-dependent in our case. This means that if we were to build a neural network to solve for optimal policy, we would need to update the network probability matrix for each backpropagation. It is computationally infeasible. Therefore, we resort to a popular heuristic combinatorial optimization method: genetic algorithm.
A.2 Evolutionary algorithm
We highlight the similarities, differences, and relevant applications of Yu et al. (2023) to our AFGA design, a fairly new paper in the evolutionary algorithm literature. While doing so, we include other relevant papers as well.
  • While evolutionary learning is popular in feature selection, task scheduling, and game theory (Yu et al., 2023; Yang et al., 2022; Li and Li, 2020), we bring a subset of evolutionary learning (GA) to a new domain: network-dependent group assignment.
  • Just like Yu et al. (2023), we treat the optimization problem as a combinatorial one in contrast to the popular DARTS method (Liu et al., 2018) which reformulates the optimization problem as a continuous one.
  • Conditionally on a random draw, the mutation stage of AFGA updates the class assignment policy “deterministically”, same as (1,1)-CELS instantiation in Yu et al. (2023).
  • Yu et al. (2023)’s (1+1)-CELS mutates parents that are hard to beat with higher probability. This concept can be adapted by our method by replacing ${\rm Multinomial}(n=1,{\rm prob}={\rm rep}(\frac{1}{\lvert C_{1}\rvert},C_{1}))$ with ${\rm Multinomial}(n=1,{\rm prob}=1/((\sum 1/\Omega z)\Omega z))$ in Algorithm 3. The modification is similar to (1+1)-CELS as we are also encouraging mutations for undesirable policy configuration (local optima in Yu et al. (2023), negatively-influencing pairs in our paper).
  • Similar to the last point, we can adapt population-based search (He and Yao, 2002; Mallipeddi and Suganthan, 2008; Tang et al., 2016; Yu et al., 2023) to our setup. At the crossover stage, we randomly select pairs of students to swap. We can assign a higher probability of being swapped to students from classrooms with low average peer effects or large variance of peer effects.
Appendix B Friendship questionnaire

### v5

@@ 4.2.1 Who are the good students/disruptive peers?
Our estimation result of (3.3) shows that if student $i$’s friends’ weighted average 6th-grade class quantile improves by 10%, his 8th-grade cognitive ability test score improves by 0.1082 on a scale of 5, or equivalently, 13% of standard deviation. To put it more concretely, the average cognitive test score is 0.3608 which lands the student at the 43rd quantile of the score distribution (the score distribution is left-skewed). An 0.1082 increase in cognitive test score improves the student’s ranking from 43rd quantile to 51st quantile. In Table 2, we show six specifications’ results. One, Linear-in-means model without classroom random effect; two, Linear-in-means model with classroom random effect; three, Instrumental variable without classroom random effect; four, Instrumental variable with classroom random effect; five, Instrumental variable with classroom-level controls; six, Instrumental variable with classroom-level controls and classroom-level random effect.
Linear-in-means IV
Peer’s Rank 0.973∗∗∗ 0.980∗∗ 1.075∗∗∗ 1.082∗∗ 0.920∗∗∗ 1.034
(0.244) (0.462) (0.269) (0.510) (0.307) (0.578)
Own Rank 1.019∗∗∗ 1.022∗∗∗ 0.995∗∗∗ 0.997∗∗∗ 0.999∗∗∗ 1.001∗∗∗
(0.038) (0.039) (0.038) (0.038) (0.038) (0.038)
Age -0.012∗∗∗ -0.012∗∗∗ -0.012∗∗∗ -0.012∗∗∗ -0.012∗∗∗ -0.012∗∗∗
(0.001) (0.001) (0.001) (0.001) (0.001) (0.001)
Sex -0.010 -0.010 -0.011 -0.010 -0.011 -0.011
(0.017) (0.017) (0.017) (0.017) (0.017) (0.017)
Father’s education 0.020∗∗∗ 0.017∗∗∗ 0.020∗∗∗ 0.018∗∗∗ 0.020∗∗∗ 0.018∗∗∗
(0.006) (0.006) (0.006) (0.006) (0.006) (0.006)
Mother’s education 0.006 0.006 0.005 0.005 0.005 0.005
(0.006) (0.006) (0.006) (0.006) (0.006) (0.006)
Ethnic nationality 0.013 0.022 0.012 0.021 0.009 0.019
(0.040) (0.040) (0.040) (0.040) (0.040) (0.040)
Teacher’s sex 0.107∗∗∗ 0.107∗∗∗
(0.034) (0.065)
Teacher’s working 0.004 0.003
experience (0.003) (0.005)
Proportion of 0.000 0.004
female (0.005) (0.010)
Proportion of 0.002 0.005
rural Hukou (0.003) (0.006)
Number of observations 5,860 5,860 5,860 5,860 5,822 5,822
Adjusted $R^{2}$ 0.382 0.382 0.384
- Log Likelihood 5,845.390 5,845.291 5,820.787
Results from first stage
Peer’s Rank 0.905∗∗∗ 0.905∗∗∗ 0.909∗∗∗ 0.909∗∗∗
(0.015) (0.015) (0.018) (0.018)
Adjusted $R^{2}$ 0.755 0.756 0.756
School FE YES YES YES YES YES YES
Class RE NO YES NO YES NO YES
Standard errors in parentheses
$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
@@
We also implement more standard methods for IV with random effect specification to estimate (3.3) (Baltagi, 1981; Balestra and Varadharajan-Krishnakumar, 1987). The estimated coefficients for peer effect are 0.975 (0.268) and 1.075 (0.270), respectively (See Table 12 in Appendix for the full results). The numbers in the parentheses are standard errors. The regression results for all six specifications and all estimation methods are very close to each other, demonstrating robustness.
The regression results opens up the strategic scope of classroom assignments. Making good students the central nodes and disruptive peers the peripheral nodes of the social network increases the entire classroom’s peer effect.
4.2.2 Examining robustness to the correlated mechanisms
We conduct robustness checks due to concerns arising from correlated mechanisms by controlling for variables related to class-level characteristics of classmates. Table 3 shows the results from regressions that include school fixed effects, as well as additional controls such as the class-level 25th and 75th percentiles of 6th-grade class rank ($z_{c}$), interaction terms between these percentiles and the proportion of female students, and an indicator for whether a student’s own rank is above the class median. We also include all baseline controls from the main regression, including class-level covariates. All results from the first to the last column indicate that peers’ ranks have a positive effect on students’ cognitive scores. However, the coefficient in the fourth column is statistically insignificant, likely due to multicollinearity between the percentile-based control variables of $z_{c}$ and the instrument classmates’ average rank (average $z_{c}$), which may inflate the standard error and reduce statistical significance.
The issue of multicollinearity becomes more pronounced when class random effects are included, as shown in columns (1) to (4) of Table 4. This is expected to some extent because all the class-level control variables and classroom indicators for the random effects are multicollinear, so the regression results rely on the normal assumption imposed on the random effect parameters. Although the estimates for peer’s rank are not statistically significant in these specifications, the direction remains the same, and the magnitude of the effects remains close to our baseline specifications.
Moreover, we conduct further robustness checks by replacing the quantile of $z_{c}$ by other measures of persistence to alleviate the multicollinearity issue. Following Golsteyn et al. (2021) and Zou (2024), persistence refers to a student’s ability to continue and persevere in their efforts, even when faced with difficulties. Persistence reflects a student’s attitude toward learning and can serve as a proxy for study-related ability. The CEPS includes questions on students’ persistence in their 6th grade. Following Zou (2024), we measure persistence using three survey items: “I would try my best to go to school even if I was not feeling very well or I had other reasons to stay at home”; “I would try my best to finish even the homework I dislike”; and “I would try my best to finish my homework even if it would take me quite a long time.” Each question is rated on a 1–5 scale (from strongly disagree to strongly agree). We sum the scores and standardize the resulting measure to have a mean of 0 and standard deviation of 1 (Zou, 2024). Then, we define students with high persistence level whose level is higher than the mean of the persistence of the sample. We calculate the proportion of students with high persistence level in each class (normal mean level and leave-out mean) and add the new control in our main regression.
The results in Table 5 show that peer rank has a significant positive effect on students’ cognitive scores. Only the estimate in column (2) is not statistically significant, but it is marginally insignificant with a p-value of 0.107. Overall, our results remain robust after controlling for persistence, further supporting our baseline results that having friends who are high-achieving in elementary school increases one’s cognitive test score in 8th grade.
(1) (2) (3) (4) (5)
Peer’s Rank 0.977 0.862 1.098∗∗ 0.684 0.980∗∗∗
(0.587) (0.458) (0.532) (0.491) (0.311)
Own Rank 0.999∗∗∗ 0.999∗∗∗ 0.999∗∗∗ 0.999∗∗∗ 0.941∗∗∗
(0.038) (0.038) (0.038) (0.038) (0.060)
25 percentile -0.038
75 percentile 0.066
25 percentile $\times$ -0.005
Proportion of female (0.012)
75 percentile $\times$ -0.012
Proportion of female (0.021)
$I[rank>medianrank]$ 0.034
Number of observations 5,822 5,822 5,822 5,822 5,822
Adjusted $R^{2}$ 0.384 0.384 0.384 0.384 0.384
Results from first stage
Peer’s Rank 0.832∗∗∗ 0.856∗∗∗ 0.834∗∗∗ 0.842∗∗∗ 0.906∗∗∗
(0.033) (0.026) (0.030) (0.028) (0.018)
Adjusted $R^{2}$ 0.757 0.757 0.757 0.757 0.756
School FE YES YES YES YES YES
Class RE NO NO NO NO NO
Standard errors in parentheses
$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
Table 3: Robustness Check with School Fixed Effects
(1) (2) (3) (4) (5)
Peer’s Rank 1.325 1.352 1.273 1.162 1.092
(1.068) (0.869) (0.968) (0.920) (0.580)
Own Rank 1.001∗∗∗ 1.001∗∗∗ 1.001∗∗∗ 1.001∗∗∗ 0.945∗∗∗
(0.038) (0.038) (0.038) (0.038) (0.059)
25 percentile -0.202
75 percentile -0.340
25 percentile $\times$ -0.007
Proportion of female (0.024)
75 percentile $\times$ -0.007
Proportion of female (0.039)
$I[rank>medianrank]$ 0.033
Number of observations 5,822 5,822 5,822 5,822 5,822
-Log Likelihood 5,820.289 5,820.118 5,823.564 5,823.092 5,822.737
Results from first stage
Peer’s Rank 0.832∗∗∗ 0.856∗∗∗ 0.834∗∗∗ 0.842∗∗∗ 0.906∗∗∗
(0.033) (0.026) (0.030) (0.028) (0.018)
Adjusted $R^{2}$ 0.757 0.757 0.757 0.757 0.756
School FE YES YES YES YES YES
Class RE YES YES YES YES YES
Standard errors in parentheses
$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
Table 4: Robustness Check with School Fixed Effects and Class Random Effects
(1) (2) (3) (4)
Peer’s Rank 0.847∗∗∗ 0.913 0.852∗∗∗ 0.962
@@ 4.2.2 Examining robustness to the correlated mechanisms
Proportion of students 0.648∗∗∗ 0.586∗∗
with high persistence level (0.143) (0.263)
Proportion of students with high 0.565∗∗∗ 0.326
persistence level (leave one out) (0.140) (0.248)
Number of observations 5,822 5,822 5,822 5,822
Adjusted $R^{2}$ 0.387 0.386
-Log Likelihood 5,818.819 5,820.449
Results from first stage
Peer’s Rank 0.907∗∗∗ 0.907∗∗∗ 0.907∗∗∗ 0.907∗∗∗
@@
Adjusted $R^{2}$ 0.757 0.757 0.757 0.757
@@
Table 5: Further Robustness Check with Persistence Level among Students
4.2.3 Who are more (or less) affected by their peers?
For completeness of our study, we examine heterogeneous peer effects of 6th-grade class rank based on gender, father’s education, and Hukou type using our instrumental variable (IV) approach. Table 6 presents the partial results. See the full results in Table 13 in the Appendix.
The first and second columns display the results by gender. Higher peers’ weighted average 6th-grade class quantile is positively associated with cognitive test scores for both males and females, though the effects are less statistically significant compared to the main regression. The effect for males is marginally significant, while the effect for females is marginally insignificant; however, the magnitudes are similar across both groups. These findings suggest that higher-achieving peers can positively influence cognitive outcomes for both genders, and there is no substantial gender difference in the strength of peer effects. For heterogeneity by father’s education, we classify students into two groups: those whose fathers have more than a middle school education (relatively high education) and those whose fathers have at most a middle school education (relatively low education). The third and fourth columns show that higher peers’ 6th-grade class quantile has a significantly positive effect on the cognitive scores of students whose fathers have relatively low education, but an insignificant effect for those with relatively high-educated fathers. This suggests that students from less advantaged family backgrounds may benefit more from exposure to higher-achieving peers, and students from more advantaged backgrounds are already well-supported and less reliant on peer influence. Finally, the fifth and sixth columns report the results by Hukou type. The estimated peer effects are significantly positive for both rural and non-rural Hukou students and are similar in magnitude to the main results. This indicates that peer effects from 6th-grade class rank do not differ between students with rural and non-rural Hukou.
Gender Father’s education Types of Hukou
Male Female Low High Rural Others
Peer’s Rank 1.000 0.946 1.255∗∗∗ 0.565 1.215∗∗ 1.143
(0.581) (0.592) (0.582) (0.687) (0.619) (0.650)
Own Rank 0.916∗∗∗ 1.117∗∗∗ 0.990∗∗∗ 1.037∗∗∗ 0.984∗∗∗ 1.008∗∗∗
(0.054) (0.053) (0.053) (0.056) (0.055) (0.053)
Number of observations 3,048 2,812 3,147 2,713 2,845 3,015
- Log Likelihood 3,244.771 2,611.928 3,283.262 2,561.519 2,953.243 2,924.823
School FE YES YES YES YES YES YES
Class RE YES YES YES YES YES YES
Standard errors in parentheses
$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
Table 6: Heterogeneous Peer Effects of Sixth Grade Rank on Cognitive Test Score
All the heterogeneous peer effect analyses are consistent with our classification of good students and disruptive peers. On average, high-achieving elementary school students become good students in middle school, influencing their friends positively; low-achieving elementary school students become disruptive peers in middle school, influencing their friends negatively.
@@ 4.3 Optimal classroom assignment policy design
4.3.1 Overall improvement with uncertainty quantification
This section shows that GA/AFGA improves the efficiency of the classroom assignment policy. We examine Step 3 with a fixed set of PeerNN parameters with the following steps. For a school, (1a) Randomly assign students to classrooms (reject the assignment if institutional constraint is violated and redraw until no violation), call it R1, compute average peer effect for R1. (1b) Randomly assign students to classrooms (reject the assignment if institutional constraint is violated and redraw until no violation), call it R2, apply GA to R2, call the output policy R2+ compute average peer effect for R2+. (1c) Repeat (1b) two more times and call the output policy R3+ and R4+. Compute the average peer effect for R3+ and R4+. (1d) Repeat step (1b) and (1c) with AFGA with the weight on equity set as 0.5, 1 and 1.5. Repeat steps (1a-1d) for all schools. For each school, we simulate 13 policies in total, 1 random assignment (R1), 4 of our methods times 3 repetitions (R2+,R3+,R4+) for each method.
Table 7 reports our methods’ improvement in average peer effect in percentage over a random assignment policy. Here are the 5 quantile (we report one-sided results because we only care about whether our policy performs better than random assignment), median, and mean of the improvement.
Method 5% Quantile Median Mean
GA 0.2413 1.9129 2.0625
AFGA 0.5 0.0477 1.8207 1.9358
AFGA 1 0.0290 1.7471 1.8242
AFGA 1.5 0.0164 1.1570 1.3175
Table 7: Performance Comparison of Methods
The improvement is positive and statistically significant at 5% level. At the median level, the improvement is 1.9% for GA and 1.2% for AFGA when $\phi=\rho=1.5$. This result shows that the heuristic optimization method, GA, makes non-trivial modifications to the initial classroom assignment policies. Moreover, the decrease in improvement down the rows makes sense because as we increase the weight, we are sacrificing efficiency for greater equity. See Figure 7 for a better visualization of the improvement’s statistical significance and trade-off between efficiency and equity.
Figure 7: The empirical distribution of the improvement demonstrates that the improvement is largely positive. As the weight put on fairness increases, the empirical distribution shifts left.
4.3.2 Inequity problem
Although there is a statistically significant improvement in the average peer effect $\tilde{\Omega}$, we encounter an inequity problem when implementing GA. In contrast, AFGA outputs policies that are more efficient than random assignment and more equitable than GA policies.
We find it clearest to explain the problem with two classrooms from one school as an example. In this section, we use classrooms 5 and 6 from CEPS as the running example. We supplement many more examples in Appendix H to show that the extreme inequity problem exists across the entire sample, not just the example that we include in the main text.
@@ 4.3.2 Inequity problem
Under the raw assignment, many students are negatively influenced by two disruptive peers as indicated by the two darker rows and columns of the Raw Q2 plot in Figure 8. GA separates many pairs of students who suffer from such large negative peer effects by ensuring that these two disruptive peers are popular among fewer students, as evidenced by the reduced number of dark grids in GA Q1 and GA Q2 relative to Raw Q1 and Raw Q2. Moreover, GA Q2 shows that the entire classroom consists mostly of positively influencing friend pairs. Unfortunately, maximizing the average peer effect inadvertently leads to severe negative peer effects for some students. A few students face extremely negative peer effects, as indicated by the very dark grids in Figure 8 GA Q2. Since these students have to disruptive peers for $q_{ij}$ to be very negative, we can infer that GA succeeds in maximizing peer effect by predicting an exclusive friendship circle among disruptive peers. In other words, these disruptive peers are “given up” by GA in pursuit of a higher average peer effect.
@@
We deploy AFGA to address this inequity issue. AF-GA Q1 and AF-GA Q2 in Figure 8 show a much more equitable educational outcome. None of the students suffer from extremely negative peer effects as in the GA Q2.
@@
If the school principal believes this particular policy designed by AFGA overvalues equity and sacrifices efficiency too much, he can decrease $\phi$ and $\rho$ for a more efficient outcome, however, the outcome will be less equitable. Note that the school principal can predict peer effect, $\tilde{\Omega}$, for all students before implementing the policy, hence he can always choose $\phi$ and $\rho$ based on his preference over efficiency versus equity.
@@ 5 Conclusion
Admittedly, implementing our framework in real life is challenging. Nevertheless, counterfactual simulation results of our framework shed light on policy recommendations. Under endogenous spillover, an efficiency-focused objective function can severely hurt the disadvantaged subpopulation. Hence, incorporating ethical considerations and equity safeguards into the policy objective should be the standard practice under such a setup.
Appendix A Friendship questionnaire